A certain CD ROM disk can store 600 megabytes of information. If an average word requires 9.0 bytes of storage, how many words can be stored on one disk? 05.4 × 109 words O 6.7 x 107 words 021x107 words 02.0×109 words
Answer #1

600 megabytes = 600*10^6 bytes

a word required 9 bytes of storage

So no of words can stored = 600*10^6 / 9 = 6.7*10^7 words

Answer = 6.7*10^7 words